Bill and Marie create posters. They can complete a poster if Bill works for 24 minutes on it and Marie works 30 minutes on it. If Marie works 36 minutes and Bill works 20 minutes they can finish a poster. How long would it take Marie to finish a poster if she works alone?
Answer by richwmiller(17219)

24/b+30/m=1
36/m+20/b=1
b = 44, m = 66
check
24/44+30/66=1,
6/11+5/11=1
check
36/66+20/44=1
6/11+5/11=1
ok